OpenStudy (anonymous):

a student council collects aluminum pop tabs to raise money to purchase a wheelchair. A company buys the pop tabs for $0.88 per kilogram. if 1267 pop tabs have a mass of one pound, how many pop tabs are needed to purchase a wheelchair worth $1500? (Hint: 1kilogram=2.2 pounds)

OpenStudy (anonymous):

1267 * 2.2 = 2787.4 tabs is 1kg. you need 1500/0.88 = 1704.54kg, so you need 1704.54 * 2787.4 = 4,751,250 tabs

OpenStudy (anonymous):

\[\frac{1500}{0.88}2.2 * 1267=4.75125*10^6 \text{ pop} \text{ tabs}\]